OpenStudy (anonymous):

Graph the system of equations using technology. What is the x-coordinate of the solution? Round to the nearest hundredth. { 2x - 11y = 4 { 3x + 5y = 14

OpenStudy (theprofessor35):

Would you like to use the substitution method?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

see i got 4.5 but im not sure if thats right?

OpenStudy (theprofessor35):

Okay, substitution is when we replace the first equation into the other. So, lets choose one equation, 2x - 11y = 4, and solve for x here. 2x - 11y = 4 +11y +11y 2x = 4 + 11y x = 5.5y + 2

OpenStudy (theprofessor35):

Now replace the value for x into that OTHER equation.

OpenStudy (theprofessor35):

3x + 5y = 14 3( 5.5y + 2 ) + 5y = 14 16.5y + 6 + 5y = 14 21.5y + 6 = 14

OpenStudy (theprofessor35):

You think you could solve for y now? So we could move on to the next step? I need to make sure you know how to solve it. :)

OpenStudy (theprofessor35):

Oh, thats fine. Let me do it here: 21.5y + 6 = 14 -6 -6 21.5y = 8 21.5y/21.5 = 8/21.5 y = 8/21.5 Now plug that into one question to solve for x 3x + 5y = 14 3x + 5(8/21.5) = 14 3x + 40/21.5 = 14 Oh my....I'm not sure about this because it stops at a rational number.....